Use the 𝑛th term formula for polygonal numbers to find the following:

a.) 32nd term of triangular number

b.) 17th term of square number

c.) 44th term of pentagonal number


1
Expert's answer
2021-03-11T06:33:17-0500

Use the 𝑛th formula for polygonal numbers to find the following:

a.) 32nd term of a triangular number

formula =TN=n(n+1)/2

whereby n=32

32(32+1)/2

=32*33/2

=1056/2

answer.=528

b.) 17th term of square number

tn=a+(n-1)d

where a=2,d=2 and n=17

t17=2+(17-1)2

2+(16)2

2+32

=34

c.) 44th term of pentagonal number

nth=n(3n-1)/2

our n =44

44(3*44-1)/2

=44(132-1)/2

5764/2

ans.=2882